OpenStudy (anonymous):

List the next three terms in the following sequence: 1, 1, 2, 3, 5, 8, ...

OpenStudy (anonymous):

Do you notice any patterns emerging?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I think the next 3 terms are 13, 21, and 34. The pattern is 1+1=2, then 2+1=3, 2+3=5..etc. Do you get it?

OpenStudy (anonymous):

yes. heres another one. What is the missing term that makes these ratios equivalent? 1.5:3, 31.5:____

OpenStudy (anonymous):

63

OpenStudy (anonymous):

Relate a vertical line test to understanding functions